# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v1

@@ 3.5. Stability thresholds on the space of geodesic rays vs divisorial valuations
Note that both the constrained inequalities appearing in the lemma are invariant under the additive action of $\mathbb{R}$ on $C^{1}(X).$ Hence, we may as well assume that $\int u\omega_{0}=0$ and set $v\coloneqq\gamma u.$ Then $\mathcal{E}(u)=-\mathcal{J}(u)$ and the equivalence thus follows directly from $\mathcal{J}(\gamma^{-1}v)=\gamma^{-2}\mathcal{J}(v).$
Now specialize to the case when $\Delta=0$ and $\phi_{0}$ is $\mathrm{SU}(2)$–invariant, so that $\boldsymbol{m}=\boldsymbol{x}$ when $X$ is identified with the unit sphere in $\mathbb{R}^{3}$ (see Example 3.10). In this case Aubin [7, Cor 2] showed that the constrained Moser–Trudinger inequality 4.5 holds for any $\gamma<2.$ Moreover, in [55] it was shown that it fails for $\gamma>2,$ by constructing an explicit destabilizing curve $v_{t}.$ However, in contrast to the Corollary above, the corresponding curve $u_{t}$ destabilizing $\mathcal{D}_{\gamma}$ does not appear to be contained in $\text{PSH}(X,\omega_{0}).$
@@
[72]Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then the constrained Moser–Trudinger inequality 4.5 holds for $\gamma\in]0,2]$ and $C_{\gamma}=0.$
We propose the following strengthening of the previous theorem:
Conjecture 4.5.
Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant and $\gamma\in]0,2].$ Then
$F_{\gamma}(\mu)\geq 0\,\,\,\forall\mu\in\mathcal{P}(X)_{0}.$
In other words, the conjecture says that $F_{\gamma}(\mu)$ is minimized on $\mathcal{P}(X)_{0}$ by the $\mathrm{SO}(3)$–invariant probability measure $\omega_{0}/V.$ When $\gamma\leq 1$ this follows from the general fact that the Mabuchi functional is minimized on the Kähler potentials of Kähler–Einstein metrics [8]. We will show below that Conjecture 4.5 holds when $\gamma\leq 3/2,$ using Theorem 4.4. Note that the conjecture is equivalent to the inequality 3.30 for the Mabuchi functional $\mathcal{M}(u)$ on $S^{2}$ subject to Tian’s orthogonality condition for $u\in\mathcal{H}(X,\omega_{0})$ for $\epsilon=1$ and $C=0.$
@@ 4.0.1. Application to Moser–Trudinger inequalities on $S^{2}$ with Tian’s constraint
for any $u\in\mathcal{H}(X,\omega_{0})$ satisfying the orthogonality condition 3.29. The conjecture was settled in [48] and, as a consequence, a new Moser–Trudinger inequality was obtained when $X=S^{2}$, saying that the inequality 4.5 holds for some $\gamma>1$ if Aubin’s constraint on $v$ is replaced by Tian’s constraint, demanding that $v$ satisfies the orthogonality condition 3.29 and $v\in\mathcal{H}(X,\omega_{0}).$ It should be emphasized that – in contrast to Aubin’s constraint – for Tian’s constraint one needs to assume that $v\in\mathcal{H}(X,\omega_{0}),$ as shown in [87]. Here we show, building on Theorem 4.4, that the Moser–Trudinger inequality on $S^{2}$ in Theorem 4.4 also holds for $\gamma\leq 2$ and $C_{\gamma}=0$ when Aubin’s constraint is replaced by Tian’s constraint (this corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.6):
@@
Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X)$ (i.e. $\int\omega_{0}=2).$ For any $v$ in $\mathcal{H}(X,\omega_{0})$ satisfying Tian’s orthogonality condition 3.29
$\frac{1}{2}\mathcal{J}(v)-\int_{X}v\omega_{0}-\log\int_{X}e^{-v}\frac{\omega_{0}}{2}\geq 0,$
i.e. the inequality is saturated when $v$ is constant.
@@
First note that it is is equivalent to show that for any $v$ in $\mathcal{H}(X,\omega_{0})$
(4.7) $\mathcal{D}_{1}(v)\geq\frac{1}{2}\mathcal{J}(v)$
if $v$ satisfies Tian’s orthogonality condition (which, in this case, is equivalent to $\int\boldsymbol{x}\omega_{v}=0;$ see Lemma 3.9 and its proof). Indeed, since both $\mathcal{D}_{1}$ and $\mathcal{J}$ are invariant under $v\mapsto v+C,$ for any $C\in\mathbb{R},$ we may as well assume that $\int v\omega_{0}=0.$ Then $\mathcal{E}(v)=\mathcal{J}(v)$ which yields the equivalence in question. Now take $u$ in $\mathcal{H}(X,\omega_{0})$ and consider the right action of $\mathcal{G}$ on $\mathcal{H}(X,\omega_{0})$ induced by the identification between $\phi\in\text{PSH}(L)$ and $u\coloneqq\phi-\phi_{0},$ where $\phi_{0}$ is a fixed metric on $K_{X}$ with (normalized) curvature form $\omega_{0}.$ In other words, $g^{*}\phi=u\cdot g+\phi_{0}$ and it follows that
@@
by applying the inequality 4.7 to $v\coloneqq u\cdot g_{0}.$ Since $\mathcal{J}(u)=E(\omega_{u}/V)$ it follows from Prop 3.4 that if $u\in\mathcal{H}(X,\omega_{0})$ and the integral of $\boldsymbol{x}\omega_{u}$ vanishes (which in the present case amounts to Tian’s constraint, by Lemma 3.9), then the inf above is attained when $g$ is the identity, showing that the inequality 4.7 indeed holds if $u$ satisfies Tian’s constraint. ∎
As a consequence, we deduce the inequality in Conjecture 4.5 for $\gamma\leq 3/2$ (which corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.6):
@@
Assume $\gamma\leq 3/2.$ Then, for all measures $\mu$ in $\mathcal{P}(X)_{0}$
@@
Expressing $\mu=\omega_{u}/V$ and combining the inequality 4.4 with the inequality 4.7 gives
@@
$\mathcal{D}_{\gamma}(v)\geq 0,\,\,\text{for}\,\gamma\in[0,2]$
if $\gamma v$ satisfies Tian’s orthogonality condition and $\gamma v\in\mathcal{H}(X,\omega_{0}).$ The former condition is equivalent to $\int\boldsymbol{x}\omega_{v}=0$ (using that, by homogeneity, $v$ satisfies Tian’s orthogonality condition, i.e. $0=\int\boldsymbol{x}v\omega_{0},$ which is proportional to $\int\boldsymbol{x}\omega_{v},$ using integration by parts). Furthermore, the condition that $\gamma v\in\mathcal{H}(X,\omega_{0})$ is equivalent to $\omega_{v}\geq(1-1/\gamma)\omega_{0},$ whose right hand side is non-negative, by assumption. We can thus apply the inequality 4.4 to $v$ to conclude that $F_{\gamma}(\frac{\omega_{v}}{V})\geq 0.$
@@ 4.1. Spontaneous symmetry breaking
In this section we take the corresponding fixed element $\phi_{0}\in\mathcal{H}(-K_{(\mathbb{P}^{1},\Delta_{w})})$ to be induced by the Fubini–Study metric, i.e. to be $\mathrm{SU}(2)$–invariant (Example 3.10) and the points $x_{0}$ and $x_{1}$ as the north and south pole of $X,$ identified with the unit sphere in $\mathbb{R}^{3}.$ Given $(X,\Delta_{w}),$ denote by $\mathcal{P}(X)_{0}^{S^{1}}$ the subspace of $\mathcal{P}(X)_{0}$ consisting of all $S^{1}$–invariant measures $\mu,$ with respect to the standard $S^{1}$–action fixing $x_{0}$ and $x_{1},$ i.e. rotation around the $x_{3}$–axis. Likewise, we denote by $C^{1}(X)_{0}^{S^{1}}$ the space of all $S^{1}$–invariant $u\in C^{1}(X)$ satisfying the moment condition 4.2.
@@
The previous proposition reveals that a phenomenon of spontaneous symmetry breaking arises when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[.$ Indeed, in this case, by Thm 3.13, Cor 4.1 and the previous proposition,
@@ 4.1.1. Failure of symmetric rearrangements
In [37, 38] and [9], the sharp logarithmic HLS inequality on the two-sphere $X$ (corresponding to the case $w=0$ and $\gamma=1)$ was shown using symmetric decreasing rearrangements. The arrangement was defined with respect to Lebesgue measure on $\mathbb{R}^{2}$ in [37, 38] (using stereographic projection), while in [9] it was defined with respect to the $\mathrm{SO}(3)$–invariant measure on the two-sphere. Both these two rearrangements define maps $\mathcal{P}(X)\rightarrow\mathcal{P}(X)^{S^{1}}$ decreasing the functional $F_{1}$ (and the arrangement in [9] decreases $F_{\gamma}$ for any $\gamma>0).$ However, neither of the rearrangements preserve the space $\mathcal{P}(X)_{0}$ (for any given weight $w).$ Moreover, the spontaneous symmetry breaking exhibited in formula 4.8 implies that when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[$ there is no map $\mathcal{P}(X)_{0}\rightarrow\mathcal{P}(X)_{0}^{S^{1}}$ that decreases the functional $F_{\gamma}.$

### v2

@@ 3.5. Stability thresholds on the space of geodesic rays vs divisorial valuations
Note that both the constrained inequalities appearing in the lemma are invariant under the additive action of $\mathbb{R}$ on $C^{1}(X).$ Hence, we may as well assume that $\int u\omega_{0}=0$ and set $\varphi\coloneqq\gamma u.$ Then $\mathcal{E}(u)=-\mathcal{J}(u)$ and the equivalence thus follows directly from $\mathcal{J}(\gamma^{-1}\varphi)=\gamma^{-2}\mathcal{J}(\varphi).$
Now specialize to the case when $\Delta=0$ and $\phi_{0}$ is $\mathrm{SU}(2)$–invariant, so that $\boldsymbol{m}=\boldsymbol{x}$ when $X$ is identified with the unit sphere in $\mathbb{R}^{3}$ (see Example 3.10). In this case Aubin [7, Cor 2] showed that the constrained Moser–Trudinger inequality 4.8 holds for any $\gamma<2.$ Moreover, in [55] it was shown that it fails for $\gamma>2,$ by constructing an explicit destabilizing curve $v_{t}.$ However, in contrast to the corollary above, the corresponding curve $u_{t}$ destabilizing $\mathcal{D}_{\gamma}$ does not appear to be contained in $\text{PSH}(X,\omega_{0}).$
@@
[72] Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then the constrained Moser–Trudinger inequality 4.8 holds for $\gamma\in]0,2]$ and $C_{\gamma}=0.$
In Theorem 5.1 we will prove strengthening of this result, saying that the minimum of the corresponding free energy $F_{\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is attained at the $SO(3)$-invariant probability measure for any $\gamma\in]0,2]$ (by the inequality 4.6 this indeed implies Theorem 4.5). But we first give some applications of Theorem 4.5 in Kähler geometry and indicate why the constrained Moser-Trudinger inequality in Theorem 4.5 is, a priori, weaker than the sharp lower bound on $F_{\gamma}$ that will be established in Theorem 5.1.
@@ 4.0.1. Application to Moser–Trudinger inequalities on $S^{2}$ with Tian’s constraint
for any $u\in\mathcal{H}(X,\omega_{0})$ satisfying the orthogonality condition 3.31. The conjecture was settled in [48] (with non-effective constants $\epsilon$ and $C)$ and, as a consequence, a new Moser–Trudinger inequality was obtained when $X=S^{2}$, saying that the inequality 4.8 holds for some $\gamma>1$ if Aubin’s constraint on $v$ is replaced by Tian’s constraint, demanding that $v$ satisfies the orthogonality condition 3.31 and $v\in\mathcal{H}(X,\omega_{0}).$ It should be emphasized that – in contrast to Aubin’s constraint – for Tian’s constraint one needs to assume that $v\in\mathcal{H}(X,\omega_{0}),$ as shown in [87]. Here we show, using Theorem 4.5, that the Moser–Trudinger inequality on $S^{2}$ in Theorem 4.5 also holds for $\gamma\leq 2$ and $C_{\gamma}=0$ when Aubin’s constraint is replaced by Tian’s constraint (this corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.9):
@@
Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X).$ For any $\varphi$ in $\mathcal{H}(X,\omega_{0})$ satisfying Tian’s orthogonality condition 3.31
$\frac{1}{2}\mathcal{J}(\varphi)-\int_{X}\varphi\frac{\omega_{0}}{2}-\log\int_{X}e^{-\varphi}\frac{\omega_{0}}{2}\geq 0,$
i.e. the inequality is saturated when $\varphi$ is constant. Equivalently,
(4.10) $\mathcal{D}_{1}(\varphi)\geq\frac{1}{2}\mathcal{J}(\varphi).$
@@
First observe that the equivalence stated in the proposition follows from the fact that both $\mathcal{D}_{1}$ and $\mathcal{J}$ are invariant under $\varphi\mapsto\varphi+C,$ for any $C\in\mathbb{R}.$ Hence, we may as well assume that $\int\varphi\omega_{0}=0.$ Then $\mathcal{E}(\varphi)=\mathcal{J}(\varphi)$ which yields the equivalence in question. Note also that $\varphi$ satisfies Tian’s orthogonality condition iff $\int\boldsymbol{x}\omega_{\varphi}=0;$ see Lemma 3.9 and its proof. Now take $u$ in $\mathcal{H}(X,\omega_{0})$ and consider the right action of $\mathcal{G}$ on $\mathcal{H}(X,\omega_{0})$ induced by the identification between $\phi\in\text{PSH}(L)$ and $u\coloneqq\phi-\phi_{0},$ where $\phi_{0}$ is a fixed metric on $K_{X}$ with (normalized) curvature form $\omega_{0}.$ In other words, $g^{*}\phi=u\cdot g+\phi_{0}$ and it follows that
@@
by applying the inequality 4.10 to $\varphi\coloneqq u\cdot g_{0}.$ Since $\mathcal{J}(u)=E(\omega_{u}/V)$ it follows from Prop 3.4 that if $u\in\mathcal{H}(X,\omega_{0})$ and the integral of $\boldsymbol{x}\omega_{u}$ vanishes (which in the present case amounts to Tian’s constraint, by Lemma 3.9), then the inf above is attained when $g$ is the identity, showing that the inequality 4.10 indeed holds if $u$ satisfies Tian’s constraint. ∎
As a consequence, we deduce the following weaker form of Theorem 5.1:
@@
Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X)$ and assume $\gamma\leq 3/2.$ Then, for all measures $\mu$ in $\mathcal{P}(X)_{0}$
@@
Expressing $\mu=\omega_{u}/V$ and combining the inequality 4.7 with the inequality 4.10 gives
@@
$\mathcal{D}_{\gamma}(\varphi)\geq 0,\,\,\text{for}\,\gamma\in[0,2]$
if $\gamma\varphi$ satisfies Tian’s orthogonality condition and $\gamma\varphi\in\mathcal{H}(X,\omega_{0}).$ The former condition is equivalent to $\int\boldsymbol{x}\omega_{\varphi}=0$ (using that, by homogeneity, $\varphi$ satisfies Tian’s orthogonality condition, i.e. $0=\int\boldsymbol{x}\varphi\omega_{0},$ which is proportional to $\int\boldsymbol{x}\omega_{\varphi},$ using integration by parts). Furthermore, the condition that $\gamma\varphi\in\mathcal{H}(X,\omega_{0})$ is equivalent to $\omega_{\varphi}\geq(1-1/\gamma)\omega_{0},$ whose right hand side is non-negative, by assumption. We can thus apply the inequality 4.7 to $\varphi$ to conclude that $F_{\gamma}(\frac{\omega_{\varphi}}{V})\geq 0.$
@@ 4.1. Spontaneous symmetry breaking
In this section we take the corresponding fixed element $\phi_{0}\in\mathcal{H}(-K_{(\mathbb{P}^{1},\Delta_{w})})$ to be $\mathrm{SU}(2)$–invariant (Example 3.10) and the points $x_{0}$ and $x_{1}$ as the north and south pole of $X,$ identified with the unit sphere $S^{2}$ in $\mathbb{R}^{3}.$ Given $(X,\Delta_{w}),$ denote by $\mathcal{P}(X)_{0}^{S^{1}}$ the subspace of $\mathcal{P}(X)_{0}$ consisting of all $S^{1}$–invariant measures $\mu,$ with respect to the standard $S^{1}$–action fixing $x_{0}$ and $x_{1},$ i.e. rotation around the $x_{3}$–axis. Likewise, we denote by $C^{1}(X)_{0}^{S^{1}}$ the space of all $S^{1}$–invariant $u\in C^{1}(X)$ satisfying the moment condition 4.5.
@@
The previous proposition reveals that a phenomenon of spontaneous symmetry breaking arises when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[.$ Indeed, in this case, by Thm 3.13, Cor 4.2 and the previous proposition,
@@ 4.1.1. Failure of symmetric rearrangements
In [37, 38] and [9], the sharp logarithmic HLS inequality on the two-sphere $X$ (corresponding to the case $w=0$ and $\gamma=1)$ was shown using symmetric decreasing rearrangements. The arrangement was defined with respect to Lebesgue measure on $\mathbb{R}^{2}$ in [37, 38] (using stereographic projection), while in [9] it was defined with respect to the $\mathrm{SO}(3)$–invariant measure on the two-sphere. Both these two rearrangements define maps $\mathcal{P}(X)\rightarrow\mathcal{P}(X)^{S^{1}}$ decreasing the functional $F_{1}$ (and the arrangement in [9] decreases $F_{\gamma}$ for any $\gamma>0).$ However, neither of the rearrangements preserve the space $\mathcal{P}(X)_{0}$ (for any given weight $w).$ Moreover, the spontaneous symmetry breaking exhibited in formula 4.11 implies that when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[$ there is no map $\mathcal{P}(X)_{0}\rightarrow\mathcal{P}(X)_{0}^{S^{1}}$ that decreases the functional $F_{\gamma}.$